OpenStudy (taytot3299):

What is the general form of the equation of a circle with center at (a, b) and radius of length m? x2 + y2 − 2ax − 2by + (a2 + b2 − m2) = 0 x2 + y2 + 2ax + 2by + (a2 + b2 − m2) = 0 x2 + y2 − 2ax − 2by + (a + b − m2) = 0 x2 + y2 + 2ax + 2by + a2 + b2 = -m2

OpenStudy (jiteshmeghwal9):

\[(x-a)^2+(y-b)^2=m^2\]

OpenStudy (firedragonrage):

That would be a circle lol just kidding idk.

OpenStudy (taytot3299):

I attached the answers it gives me

OpenStudy (jiteshmeghwal9):

\[(x-a)^2+(y-b)^2=m^2\]\[x^2-2ax+a^2+y^2-2by+b^2=m^2\]simplify it further

OpenStudy (jackthegreatest):

\[x^2 + y^2 -2ax -2by + (a^2+b^2-m^2)\]

OpenStudy (jackthegreatest):

= 0

OpenStudy (mathmale):

It'd be helpful if you were to write out the standard equation of a circle with center at (h,k) and radius r (or radius m) and then compare this to each of the four given equations.